Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-11-26

Observers: Z.Kostrzewa-Rutkowska, K.Lopez
TAT: J. Chen, M. Fraser, J. Sollerman
Support team: P. Jonker, G. Cannizzaro, F. Onori
Instruments: EFOSC2

Weather conditions: 
evening UT 23:45: thin clouds, humidity 47%
morning UT 8:30: clear, humidity 28%

Calibrations:
EFOSC2 flats, bias, arcs daytime, arcs morning

LOG:

UT      Target           Type       Filter/Grism ExpTime[s]  Slit   Seeing  Airmass    Comments
00:06   Feige110         STD        gr11/gr13/gr16          1.0/1.5 1.09     1.10
00:28   AT2017hya        Followup   gr11/gr16    1800/1800  1.0     1.20     1.02      image quality issues, aborted
00:48   AT2017hya        Followup   gr11/gr16    1800/1800  1.0     0.95     1.05/1.09 restarted
02:08   AT2017gpp        Followup   gr13         2700       1.0     0.84     1.48
03:03   AT2017hcc        Followup   gr11/gr16    600/600    1.0     0.96     1.35/1.40
03:45   ATLAS17nic       Class      gr13         900        1.0     1.50     1.16      
04:10                                                                                  image quality fixing
04:20   AT2017iit        Class      gr13         600        1.0     1.02     1.54
04:40   L745a            STD        gr11/gr13/gr16          1.0/1.5 1.33     1.44
05:05   MASTERJ034412..  Followup   gr11         3600       1.0     1.00     1.04
06:15   ATLAS17mwz       Followup   gr11/gr16    1500/1500  1.0     0.77     1.11/1.15
07:22   ATLAS17nkf       Class      gr13         900        1.0     0.80     1.21
07:48   ATLAS17nfm       Class      gr13         600        1.0     0.82     1.58
08:15   L745a            STD        gr11/gr13/gr16          1.0/1.5 0.82     1.03

Comments